9. LounchPad. An economy has 100 people divided among the folteoving groups: 25 have full-time jobs, 20 hare one part-time job, 5 have two part-time jobs, 10 would like to work and are looking for jobs, 10 would like to work but are so discouraged they have given up looking, 10 are running their own businesses, 10 are retired, and 10 are small children. a. Calculate the labor force and the labor-force participation rate unemployment rate. as measured by the household survey and as b. Calculate the number of unemployed and the c. Calculate total employment in two ways: measured by the establishment survey.
Answer #1

Answer 9 a) Labour Force = Employed + Unemployed

Labour Force = 60+ 10= 70

Discouraged workers= 10

Retired person= 10

Small children= 10

Labour Force participation rate=( Labour Force/ Adult population)*100= ( 70/80)*100= 87.5%

b: Number of Unemployed person = 10

Labour Force = 70

Unemployment rate = (Number of Unemployed/ Labour Force)*100= ( 10/70)*100= 14.28%

c . Total unemployment as measure by household survey : In this method , the person who are actively searching for a job is considered to be unemployed. It means that person who is discouraged are not taken into an account.

Here, Unemployed labour= 10

Labour Force= 70

Unemployment rate= (10/70)*100= 14.28%

Where as EstImated survey method. People who claim the benefit of unemployed is part of unemployment rate

Unemployed = 20

Unemployment rate= (20/80)*100= 25%

Estimated unemployment is always higher than the household survey method.